OpenStudy (anonymous):

Sketch a directed line segment in R^2 representing each of the following vectors. A. X1=[ -2] [ 3]

OpenStudy (anonymous):

First draw a coordinate system in \(\mathbb R^2\) (might as well be rectangular)...

OpenStudy (anonymous):



OpenStudy (anonymous):

the vector you are given I am reading as\[\vec x_1=\langle-2,3\rangle\]which is a vector that point from the origin to the point \((-2,3)\) so draw that in. make sure the point of the arrow is on the point specified.

OpenStudy (anonymous):

um, no it said on the worksheet x1= [-2] [ 3]. The bracket covers the entire -2 and 3.

OpenStudy (anonymous):

the way you write it makes very little difference

OpenStudy (anonymous):

\[\vec x_1=\left[\begin{array}1-2\\3\end{array}\right]=\langle-2,3\rangle=-2\hat i+3\hat j\]all equivalent

OpenStudy (anonymous):

okay